OpenStudy (anonymous):

Kamal can do a piece of work in 12 days which Debashis can finish in 10 days. If they work at it on alternate days with Debashis begining, in how many days, the work will be finished? a) 10 2/3 b) 10 5/6 c) 11 3/4 d) 11 2/7

OpenStudy (kropot72):

After 10 days, Debashis and Kamal will have each worked for 5 days. The amount of the work that has been done after 10 days is given by: \[\large \frac{5}{10}+\frac{5}{12}=\frac{11}{12}\] It is the turn of Debashis on the eleventh day, when he has 1/12 of the work remaining to finish. It will take Debashis 5/6 of the eleventh day to finish the work completely. Therefore the total time is 10 5/6 days.

OpenStudy (kropot72):

@dhiwagm Are you there?